### v1

# Adaptive Batch Size for Privately Finding Second-Order Stationary Points

Abstract
There is a gap between finding a first-order stationary point (FOSP) and a second-order stationary point (SOSP) under differential privacy constraints, and it remains unclear whether privately finding an SOSP is more challenging than finding an FOSP. Specifically, Ganesh et al. (2023) demonstrated that an $\alpha$-SOSP can be found with $\alpha=\tilde{O}(\frac{1}{n^{1/3}}+(\frac{\sqrt{d}}{n\varepsilon})^{3/7})$, where $n$ is the dataset size, $d$ is the dimension, and $\varepsilon$ is the differential privacy parameter. Building on the SpiderBoost algorithm framework, we propose a new approach that uses adaptive batch sizes and incorporates the binary tree mechanism. Our method improves the results for privately finding an SOSP, achieving $\alpha=\tilde{O}(\frac{1}{n^{1/3}}+(\frac{\sqrt{d}}{n\varepsilon})^{1/2})$. This improved bound matches the state-of-the-art for finding an FOSP, suggesting that privately finding an SOSP may be achievable at no additional cost.
1 Introduction
Privacy concerns have gained increasing attention with the rapid development of artificial intelligence and modern machine learning, particularly the widespread success of large language models. Differential privacy (DP) has become the standard notion of privacy in machine learning since it was introduced by Dwork et al. (2006). Given two neighboring datasets, $\mathcal{D}$ and $\mathcal{D}^{\prime}$, differing by a single item, a mechanism $\mathcal{M}$ is said to be $(\varepsilon,\delta)$-differentially private if, for any event $\mathcal{X}$, it holds that:
$\Pr[\mathcal{M}(\mathcal{D})\in\mathcal{X}]\leq e^{\varepsilon}\Pr[\mathcal{M}(\mathcal{D}^{\prime})\in\mathcal{X}]+\delta.$
In this work, we focus on the stochastic optimization problem under the constraint of DP. The loss function is defined below:
$F_{\mathcal{P}}(x):=\operatorname*{\mathbb{E}}_{z\sim\mathcal{P}}f(x;z),$
where the functions may be non-convex, the underlying distribution $\mathcal{P}$ is unknown, and we are given a dataset $\mathcal{D}=\{z_{i}\}_{i\in[n]}$ drawn i.i.d. from $\mathcal{P}$. Notably, our goal is to design a private algorithm with provable utility guarantees under the i.i.d. assumption.
Minimizing non-convex functions is generally challenging and often intractable, but most models used in practice are not guaranteed to be convex. How, then, can we explain the success of optimization methods in practice? One possible explanation is the effectiveness of Stochastic Gradient Descent (SGD), which is well-known to be able to find an $\alpha$-first-order stationary point (FOSP) of a non-convex function $f$—that is, a point $x$ such that $\|\nabla f(x)\|\leq\alpha$—within $O(1/\alpha^{2})$ steps (Nesterov, 1998). However, FOSPs can include saddle points or even local maxima. Thus, we focus on finding second-order stationary points (SOSP), for the non-convex function $F_{\mathcal{P}}$.
Non-convex optimization has been extensively studied in recent years due to its central role in modern machine learning, and we now have a solid understanding of the complexity involved in finding FOSPs and SOSPs (Ghadimi and Lan, 2013; Agarwal et al., 2017; Carmon et al., 2020; Zhang et al., 2020). Variance reduction techniques have been shown to improve the theoretical complexity, leading to the development of several promising algorithms such as Spider (Fang et al., 2018), SARAH (Nguyen et al., 2017), and SpiderBoost (Wang et al., 2019c). More recently, private non-convex optimization has emerged as an active area of research (Wang et al., 2019b; Arora et al., 2023; Gao and Wright, 2023; Ganesh et al., 2023; Lowy et al., 2024; Menart et al., 2024).
1.1 Our Main Result
In this work, we study how to find the SOSP of $F_{\mathcal{P}}$ privately. Let us formally define the FOSP and the SOSP. For more on Hessian Lipschitz continuity and related background, see the preliminaries in Section 2.
Definition 1.1 (FOSP).
For $\alpha\geq 0$, we say a point $x$ is an $\alpha$-first-order stationary point ($\alpha$-FOSP) of a function $g$ if $\|\nabla g(x)\|\leq\alpha$.
Definition 1.2 (SOSP, Nesterov and Polyak (2006); Agarwal et al. (2017)).
For a function $g:\mathbb{R}^{d}\to\mathbb{R}$ which is $\rho$-Hessian Lipschitz, we say a point $x\in\mathbb{R}^{d}$ is $\alpha$-second-order stationary point ($\alpha$-SOSP) of $g$ if $\|\nabla g(x)\|_{2}\leq\alpha\;\bigwedge\;\nabla^{2}g(x)\succeq-\sqrt{\rho\alpha}I_{d}$.
Given the dataset size of $n$, privacy parameters $\varepsilon,\delta$, and functions defined over $d$-dimension space, Ganesh et al. (2023) proposed a private algorithm that can find an $\alpha_{S}$-SOSP for $F_{\mathcal{P}}$ with
$\alpha_{S}=\tilde{O}\big{(}\frac{1}{n^{1/3}}+(\frac{\sqrt{d}}{n\varepsilon})^{3/7}\big{)}.$
However, as shown in Arora et al. (2023), the state-of-the-art bound for privately finding an $\alpha_{F}$-FOSP is tighter:  [^1] As proposed by Lowy et al. (2024), allowing exponential running time enables the use of the exponential mechanism to find a warm start, which can further improve the bounds for both FOSP and SOSP.
$\alpha_{F}=\tilde{O}\big{(}\frac{1}{n^{1/3}}+(\frac{\sqrt{d}}{n\varepsilon})^{1/2}\big{)}$
When the privacy parameter $\varepsilon$ is sufficiently small, we observe that $\alpha_{F}\ll\alpha_{S}$. This raises the question: is finding an SOSP under differential privacy constraints more difficult than finding an FOSP?
This work improves the results of Ganesh et al. (2023). Specifically, we present an algorithm that finds an $\alpha$-SOSP with privacy guarantees, where:
$\alpha=\tilde{O}(\alpha_{F})=\tilde{O}\big{(}\frac{1}{n^{1/3}}+(\frac{\sqrt{d}}{n\varepsilon})^{1/2}\big{)}.$
This improved bound suggests that when we try to find the stationary point privately, we can find the SOSP for free under additional (standard) assumptions.
It is also worth noting that, our improvement primarily affects terms dependent on the privacy parameters. In the non-private setting, as $\varepsilon\to\infty$ (i.e., without privacy constraints), all the results discussed above achieve a bound of $\tilde{O}(1/n^{1/3})$, which matches the non-private lower bound established by Arjevani et al. (2023) in high-dimensional settings (where $d\geq\tilde{\Omega}(1/\alpha^{4})$). However, to our knowledge, whether this non-private term of $\tilde{O}(1/n^{1/3})$ can be further improved in low dimension remains an open question.
1.2 Overview of Techniques
In this work, we build on the SpiderBoost algorithm framework, similar to prior approaches (Arora et al., 2023; Ganesh et al., 2023), to find second-order stationary points (SOSP) privately. At a high level, our method leverages two types of gradient oracles: $\mathcal{O}_{1}(x)\approx\nabla f(x)$, which estimates the gradient at point $x$, and $\mathcal{O}_{2}(x,y)\approx\nabla f(x)-\nabla f(y)$, which estimates the gradient difference between two points, $x$ and $y$. When performing gradient descent, to compute the gradient estimator $\nabla_{t}$ at point $x_{t}$, we can either use $\nabla_{t}=\mathcal{O}_{1}(x_{t})$ for a direct estimate, or $\nabla_{t}=\nabla_{t-1}+\mathcal{O}_{2}(x_{t},x_{t-1})$ to update based on the previous gradient. In our setting, $\mathcal{O}_{1}$ is more accurate but incurs higher computational or privacy costs.
The approach in Arora et al. (2023) adopts SpiderBoost by querying $\mathcal{O}_{1}$ periodically: they call $\mathcal{O}_{1}$ once and then use $\mathcal{O}_{2}$ for $q$ subsequent queries, controlled by a hyperparameter $q$. Their method ensures the gradient estimators are sufficiently accurate on average, which works well for finding first-order stationary points (FOSP). However, finding an SOSP, where greater precision is required, presents additional challenges when relying on average-accurate gradient estimators.
To address this, Ganesh et al. (2023) introduced a variable called $\mathrm{drift}_{t}:=\sum_{i=t_{0}+1}^{t}\|x_{i}-x_{i-1}\|^{2}$, where $t_{0}$ is the index of the last iteration when $\mathcal{O}_{1}$ was queried. If $\mathrm{drift}_{t}$ remains small, the gradient estimator stays accurate enough, allowing further queries of $\mathcal{O}_{2}$. However, if $\mathrm{drift}_{t}$ grows large, the gradient estimator’s accuracy deteriorates, signaling the need to query $\mathcal{O}_{1}$ for a fresh, more accurate estimate. This modification enables the algorithm to maintain the precision necessary for privately finding an SOSP.
Our improvement introduces two new components: the use of the tree mechanism instead of using the Gaussian mechanism as in Ganesh et al. (2023), and the implementation of adaptive batch sizes for constructing $\mathcal{O}_{2}$.
In the prior approach using the Gaussian mechanism, a noisy gradient estimator $\nabla_{t-1}$ is computed, and the next estimator is updated via $\nabla_{t}=\nabla_{t-1}+\mathcal{O}_{2}(x_{t},x_{t-1})+g_{t}$, where $g_{t}$ is Gaussian noise added to preserve privacy. Over multiple iterations, the accumulation of noise $\sum g_{t}$ can severely degrade the accuracy of the gradient estimator, requiring frequent re-queries of $\mathcal{O}_{1}$. On the other hand, the tree mechanism mitigates this issue when frequent queries to $\mathcal{O}_{2}$ are needed.
However, simply replacing the Gaussian mechanism with the tree mechanism and using a fixed batch size does not yield optimal results. In worst-case scenarios, where the function’s gradients are large, the $\mathrm{drift}$ grows quickly, necessitating frequent calls to $\mathcal{O}_{1}$, which diminishes the advantages of the tree mechanism.
To address this, we introduce adaptive batch sizes. In Ganesh et al. (2023), the oracle $\mathcal{O}_{2}$ is constructed by drawing a fixed batch of size $B$ from the unused dataset and outputting $\mathcal{O}_{2}(x_{t},x_{t-1}):=\sum_{z\in S_{t}}\frac{\nabla f(x_{t};z)-\nabla f(x_{t-1};z)}{B}$. Given an upper bound on $\mathrm{drift}$, they guaranteed that $\|x_{t}-x_{t-1}\|\leq D$ for some parameter $D$, thereby bounding the sensitivity of $\mathcal{O}_{2}$.
In contrast, we dynamically adjust the batch size in proportion to $\|x_{t}-x_{t-1}\|$, setting $B_{t}\propto\|x_{t}-x_{t-1}\|$, and compute $\mathcal{O}_{2}(x_{t},x_{t-1}):=\sum_{z\in S_{t}}\frac{\nabla f(x_{t};z)-\nabla f(x_{t-1};z)}{B_{t}}$. Fixed batch sizes present two drawbacks: (i) when $\|x_{t}-x_{t-1}\|$ is large, the gradient estimator has higher sensitivity and variance, leading to worse estimate accuracy; (ii) when $\|x_{t}-x_{t-1}\|$ is small, progress in terms of function value decrease is limited. Using a fixed batch size forces us to handle both cases simultaneously: we must add noise and analyze accuracy assuming a worst-case large $\|x_{t}-x_{t-1}\|$, but for utility analysis, we pretend $\|x_{t}-x_{t-1}\|$ is small to examine the function value decrease. The adaptive batch size resolves this paradox: it allows us to control sensitivity and variance adaptively. When $\|x_{t}-x_{t-1}\|$ is small, we decrease the batch size but can still control the variance and sensitivity; when it is small, the function value decreases significantly, aiding in finding an SOSP.
By combining the tree mechanism with adaptive batch sizes, we improve the accuracy of gradient estimation and achieve better results for privately finding an SOSP.
1.3 Other Related Work
A significant body of literature on private optimization focuses on the convex setting, where it is typically assumed that each function $f(;z)$ is convex for any $z$ in the universe (e.g., (Chaudhuri et al., 2011; Bassily et al., 2014, 2019; Feldman et al., 2020; Asi et al., 2021; Kulkarni et al., 2021; Carmon et al., 2023)).
The tree mechanism, originally introduced by the differential privacy (DP) community (Dwork et al., 2010; Chan et al., 2011) for the continual observation, has inspired tree-structure private optimization algorithms like Asi et al. (2021); Bassily et al. (2021); Arora et al. (2023); Zhang et al. (2024). Some prior works have explored adaptive batch size techniques in optimization. For instance, De et al. (2016) introduced adaptive batch sizing for stochastic gradient descent (SGD), while Ji et al. (2020) combined adaptive batch sizing with variance reduction techniques to modify SVRG and Spider algorithms. However, these works’ motivations and approaches to setting adaptive batch sizes differ from ours. To the best of our knowledge, we are the first to propose using adaptive batch sizes in the context of optimization under differential privacy constraints.
Most of the non-convex optimization literature assumes that the functions being optimized are smooth. Recent work has begun addressing non-smooth, non-convex functions as well, as seen in Zhang et al. (2020); Kornowski and Shamir (2021); Davis et al. (2022); Jordan et al. (2023).
2 Preliminaries
[…]
Definition 2.1 (Lipschitz, Smoothness and Hessian Lipschitz).
Let $\mathcal{K}\subseteq\mathbb{R}^{d}$. Given a twice differentiable function $f:\mathcal{K}\to\mathbb{R}$, we say $f$ is $G$-Lipschitz, if for all $x_{1},x_{2}\in\mathcal{K}$, $|f(x_{1})-f(x_{2})|\leq G\|x_{1}-x_{2}\|$; we say $f$ is $M$-smooth, if for all $x_{1},x_{2}\in\mathcal{K}$, $\|\nabla f(x_{1})-\nabla f(x_{2})\|\leq M\|x_{1}-x_{2}\|$, and we say the function $f$ is $\rho$-Hessian Lipschitz, if for all $x_{1},x_{2}\in\mathcal{K}$, we have $\|\nabla^{2}f(x_{1})-\nabla^{2}f(x_{2})\|\leq\rho\|x_{1}-x_{2}\|.$
2.1 Other Techniques
As mentioned in the introduction, we use the tree mechanism (Algorithm 1) to privatize the algorithm, whose formal guarantee is stated below:
Theorem 2.2 (Tree Mechanism, Dwork et al. (2010); Chan et al. (2011)).
Let $\mathcal{Z}_{1},\cdots,\mathcal{Z}_{\Sigma}$ be dataset spaces, and $\mathcal{X}$ be the state space. Let $\mathcal{M}_{i}:\mathcal{X}^{i-1}\times\mathcal{Z}_{i}\to\mathcal{X}$ be a sequence of algorithms for $i\in[\Sigma]$. Let $\mathcal{A}:\mathcal{Z}^{(1:\Sigma)}\to\mathcal{X}^{\Sigma}$ be the algorithm that given a dataset $Z_{1:\Sigma}\in\mathcal{Z}^{(1:\Sigma)}$, sequentially computes $X_{i}=\sum_{j=1}^{i}\mathcal{M}_{i}(X_{1:j-1},Z_{i})+\mathrm{TREE}(i)$ for $i\in[\Sigma]$, and then outputs $X_{1:\Sigma}$.
Suppose for all $i\in[\Sigma]$, and neighboring $Z_{1:\Sigma},Z_{1:\Sigma}^{\prime}\in\mathcal{Z}^{(1:\Sigma)},\|\mathcal{M}_{i}(X_{1:i-1},Z_{i})-\mathcal{M}_{i}(X_{1:i-1},Z_{i}^{\prime})\|\leq s$ for all auxiliary inputs $X_{1:i-1}\in\mathcal{X}^{i-1}$. Then setting $\sigma=\frac{4s\sqrt{\log\Sigma\log(1/\delta)}}{\varepsilon}$, Algorithm 1 is $(\varepsilon,\delta)$-DP. Furthermore, with probability at least $1-\Sigma\cdot\iota$, for all $t\in[\Sigma]:\|\mathrm{TREE}(t)\|\lesssim\sqrt{d\log(1/\iota)}\sigma$.
We also need the concentration inequality for norm-subGaussian random vectors.
Definition 2.3 (SubGaussian, and Norm-SubGaussian).
We say a random vector $x\in\mathbb{R}^{d}$ is SubGaussian ($\mathrm{SG}(\zeta)$) if there exists a positive constant $\zeta$ such that $\operatorname*{\mathbb{E}}e^{\langle v,x-\operatorname*{\mathbb{E}}x\rangle}\leq e^{\|v\|^{2}\zeta^{2}/2},\;\;\forall v\in\mathbb{R}^{d}$. We say $x\in\mathbb{R}^{d}$ is norm-SubGaussian ($\mathrm{nSG}(\zeta)$) if there exists $\zeta$ such that $\Pr[\|x-\operatorname*{\mathbb{E}}x\|\geq t]\leq 2e^{-\frac{t^{2}}{2\zeta^{2}}},\forall t\in\mathbb{R}$.
Lemma 2.4 (Hoeffding type inequality for norm-subGaussian, Jin et al. (2019)).
Let $x_{1},\cdots,x_{k}\in\mathbb{R}^{d}$ be random vectors, and for each $i\in[k]$, $x_{i}\mid\mathcal{F}_{i-1}$ is zero-mean $\mathrm{nSG}(\zeta_{i})$ where $\mathcal{F}_{i}$ is the corresponding filtration. Then there exists an absolute constant $c$ such that for any $\delta>0$, with probability at least $1-\omega$, $\|\sum_{i=1}^{k}x_{i}\|\leq c\cdot\sqrt{\sum_{i=1}^{k}\zeta_{i}^{2}\log(2d/\omega)}$, which means $\sum_{i=1}^{k}x_{i}$ is $\mathrm{nSG}(\sqrt{c\log(d)\,\sum_{i=1}^{k}\zeta_{i}^{2}})$.
Theorem 2.5 (Matrix Bernstein Inequality, Tropp et al. (2015)).
Consider a finite sequence of independent, random matrices $X_{1},\cdots,X_{n}$ with common dimensions $d\times d$ and $\operatorname*{\mathbb{E}}[X_{i}]=0,\|X_{i}\|\leq L$ a.s., $\forall i$. Let $\sigma^{2}=\|\sum_{i=1}^{n}\operatorname*{\mathbb{E}}[X_{i}^{2}]\|$. Let $S=\sum_{i=1}^{n}X_{i}$. Then for any $t\geq 0$, we have
$\Pr[\|S\|\geq t]\leq d\cdot\exp(-\frac{-t^{2}}{\sigma^{2}+Lt/3}).$
1:Input: Noise parameter $\sigma$, sequence length $\Sigma$
2:Define $\mathcal{T}:=\{(u,v):u=j\cdot 2^{\ell-1}+1,v=(j+1)\cdot 2^{\ell-1},1\leq\ell\leq\log\Sigma,0\leq j\leq\Sigma/2^{\ell-1}-1\}$
3:Sample and store $\zeta_{(u,v)}\sim\mathcal{N}(0,\sigma^{2})$ for each $(u,v)\in\mathcal{T}$
4:for $t=1,\cdots,\Sigma$ do
5:     Let $\mathrm{TREE}(t)\leftarrow\sum_{(u,v)\in\mathrm{NODE}(t)}\zeta_{(u,v)}$
6:end for
7:Return: $\mathrm{TREE}(t)$ for each $t\in[\Sigma]$
8:
9:Function NODE:
10:Input: index $t\in[\Sigma]$
11:Initialize $S=\{\}$ and $k=0$
12:for $i=1,\cdots,\lceil\log\Sigma\rceil$ while $k<t$ do
13:     Set $k^{\prime}=k+2^{\lceil\log\Sigma\rceil-i}$
14:     if $k^{\prime}\leq k$ then
15:         $S\leftarrow S\cup\{(k+1,k^{\prime})\}$, $k\leftarrow k^{\prime}$
16:     end if
17:end for
Algorithm 1 Tree Mechanism
3 SOSP
We make the following assumption for our main result.
Assumption 3.1.
Let $G,\rho,M,B>0$. Any function drawn from $\mathcal{P}$ is $G$-Lipschitz, $\rho$-Hessian Lipschitz, and $M$-smooth, almost surely. Moreover, we are given a public initial point $x_{0}$ such that $F_{\mathcal{P}}(x_{0})-\inf_{x}F_{\mathcal{P}}(x)\leq B$.
We modify the Stochastic SpiderBoost used in Ganesh et al. (2023) and state it in Algorithm 2. The following standard lemma plays a crucial role in finding stationary points of smooth functions:
Lemma 3.2.
Assume $F$ is $M$-smooth and let $\eta=1/M$. Let $x_{t+1}=x_{t}-\eta\widetilde{\nabla}$. Then we have $F(x_{t+1})\leq F(x_{t})+\eta\|\widetilde{\nabla}_{t}\|\cdot\|\nabla F(x_{t})-\widetilde{\nabla}_{t}\|-\frac{\eta}{2}\|\widetilde{\nabla}_{t}\|^{2}.$ Moreover, if $\|\nabla F(x_{t})\|\geq\gamma$ and ~$\|\widetilde{\nabla}_{t}-\nabla F(x_{t})\|\leq\gamma/4$, we have
$F(x_{t+1})-F(x_{t})\leq-\eta\|\widetilde{\nabla}_{t}\|^{2}/16.$
Proof.
By the assumption of the smoothness, we know
$F(x_{t+1})$ $\leq F(x_{t})+\langle\nabla F(x_{t}),x_{t+1}-x_{t}\rangle+\frac{M}{2}\|x_{t+1}-x_{t}\|^{2}$
$=F(x_{t})-\eta/2\|\widetilde{\nabla}_{t}\|^{2}-\langle\nabla F(x_{t})-\widetilde{\nabla}_{t},\eta\widetilde{\nabla}_{t}\rangle$
$\leq F(x_{t})+\eta\|\nabla F(x_{t})-\widetilde{\nabla}_{t}\|\cdot\|\widetilde{\nabla}_{t}\|-\frac{\eta}{2}\|\widetilde{\nabla}_{t}\|^{2}.$
When $\|\nabla F(x_{t})\|\geq\gamma$ and ~$\|\widetilde{\nabla}_{t}-\nabla F(x_{t})\|\leq\gamma/4$, the conclusion follows from the calculation.
One of the main challenges to finding the SOSP, compared with finding the FOSP, is showing the algorithm can escape from saddle points, which was addressed by previous results (e.g.,Jin et al. (2017)) and was established in the DP literature as well:
Lemma 3.3 (Essentially from Wang et al. (2019a)).
Under Assumption 3.1, run SGD iterations $x_{t+1}=x_{t}-\eta\nabla_{t}$, with step size $\eta=1/M$. Suppose $x_{0}$ is a saddle point satisfying $\|\nabla F(x_{0})\|\leq\alpha$ and $\textup{smin}(\nabla^{2}F(x_{0}))\leq-\sqrt{\rho\alpha}$, $\alpha=\gamma\log^{3}(dBM/\rho\iota)$. If $\nabla_{0}=\nabla F(x_{0})+\zeta_{1}+\zeta_{2}$ where $\|\zeta_{1}\|\leq\gamma$, $\zeta_{2}\sim\mathcal{N}(0,\frac{\gamma^{2}}{d\log(d/\iota)}I_{d})$, and $\|\nabla_{t}-\nabla F(x_{t})\|\leq\gamma$ for all $t\in[\Gamma]$, with probability at least $1-\iota$, one has $F(x_{\Gamma})-F(x_{0})\leq-\Omega\big{(}\frac{\gamma^{3/2}}{\sqrt{\rho}\log^{3}(\frac{dMB}{\rho\gamma\iota})}\big{)},$ where $\Gamma=\frac{M\log(\frac{dMB}{\rho\gamma\iota})}{\sqrt{\rho\gamma}}$.
Lemma 3.3 suggests that, if we meet a saddle point, then after the following $\tilde{O}(1/\sqrt{\gamma})$ steps, the function value will decrease by at least $\tilde{\Omega}(\gamma^{3/2})$. This means the function value decreases by $\tilde{\Omega}(\gamma^{2})$ on average for each step. Similar to the proof in Ganesh et al. (2023), we have the following guarantee of Stochastic SpiderBoost:
Proposition 3.4.
Under Assumption 3.1 and with gradient oracles such that ~$\|\widetilde{\nabla}_{t}-\nabla F(x_{t})\|\leq\gamma$ for any $t\in[T]$, setting $T=\tilde{O}(B/\eta\gamma^{2})$ and $\zeta=\gamma/\sqrt{\log(d/\iota)}$ and supposing it does not halt before completing all $T$ iterations, with probability at least $1-T\iota$, at least one point in the output set $\{x_{1},\cdots,x_{T}\}$ of Algorithm 2 is $\tilde{O}(\gamma)$-SOSP.
The proof intuition of Proposition 3.4 is, if we do not find an $O(\gamma)$-SOSP, then on average, the function value will at least decrease by $\Omega(\eta/\gamma^{2})$. As we know $F_{\mathcal{P}}(x_{0})\leq F_{\mathcal{P}}^{*}+B$, hence $O(B/\eta\gamma^{2})$ steps can ensure we find an $O(\gamma)$-SOSP. See more discussion on Proposition 3.4 in the Appendix.
Algorithm 2 Stochastic Spider
1:Input: Dataset $\mathcal{D}$, privacy parameters $\varepsilon,\delta$, parameters of objective function $B,M,G,\rho$, parameter $\kappa$, failure probability $\omega$, batch size parameter $b$, noise parameter $\zeta$
2:set $\mathrm{drift}_{0}=\kappa,\mathrm{frozen}_{-1}=1,\Delta_{-1}=0,\mathcal{D}_{r}\leftarrow\mathcal{D},t=0$
3:while $t<T$ and the number of unused functions is larger than $b$ do
4:     if $\mathrm{drift}_{t}\geq\kappa$ then
5:         $\nabla_{t}=\mathcal{O}_{1}(x_{t})$, $\mathrm{drift}_{t}=0$, $\mathrm{frozen}_{t}=\mathrm{frozen}_{t-1}-1$
6:     else
7:         $\Delta_{t}=\mathcal{O}_{2}(x_{t},x_{t-1})$, $\nabla_{t}=\nabla_{t-1}+\Delta_{t}$, $\mathrm{frozen}_{t}=\mathrm{frozen}_{t-1}-1$
8:     end if
9:     if ~$\|\widetilde{\nabla}_{t-1}\|\leq\gamma\log^{3}(BMd/\rho\omega)\bigwedge\mathrm{frozen}_{t-1}\leq 0$ then
10:         Set $\mathrm{frozen}_{t}=\Gamma,\mathrm{drift}_{t}=0$
11:         Set ~$\widetilde{\nabla}_{t}=\nabla_{t}+\mathrm{TREE}(t)+g_{t}$, where $g_{t}\sim\mathcal{N}(0,\frac{\zeta^{2}}{d}I_{d})$
12:     else
13:         Set ~$\widetilde{\nabla}_{t}=\nabla_{t}+\mathrm{TREE}(t)$
14:     end if
15:     $x_{t+1}=x_{t}-\eta\widetilde{\nabla}_{t},\mathrm{drift}_{t+1}=\mathrm{drift}_{t}+\|\widetilde{\nabla}_{t}\|_{2}$, $t=t+1$
16:end while
17:Return: $\{x_{1},\cdots,x_{t}\}$
Algorithm 2 follows the SpiderBoost framework. We either query $\mathcal{O}_{1}$ to estimate the gradient itself or query $\mathcal{O}_{2}$ to estimate the gradient difference over the last consecutive points. The term $\mathrm{drift}$ controls the estimated error. When $\mathrm{drift}_{t}$ is small, we know $\Delta_{t}$ is still a good estimator, and when $\mathrm{drift}_{t}$ is large, we draw a fresh estimator from $\mathcal{O}_{1}$. The term $\mathrm{frozen}$ is used for the technical purpose of applying Lemma 3.3. When we meet a potential saddle point, we add Gaussian noise $g_{t}$ to escape from the saddle point and set $\mathrm{frozen}$ to be $\Gamma$; this ensures that we won’t add the Gaussian again in the following $\Gamma$ steps.
3.1 Constructing Private Gradient Oracles
We construct the gradient oracles below in Algorithm 3.
Algorithm 3 gradient oracles
1:gradient $\mathcal{O}_{1}$
2:inputs: $x_{t}$
3:draw a batch size of $b$ among unused functions
4:return: $\frac{1}{b}\sum_{z}\nabla f(x_{t};z)$
1:gradient $\mathcal{O}_{2}$
2:inputs: $x_{t},x_{t-1}$
3:draw a batch size of $b_{t}$ among unused functions
4:return: $\frac{1}{b_{t}}\sum_{z}(\nabla f(x_{t};z)-\nabla f(x_{t-1};z))$
Lemma 3.5 (Gradient oracles with bounded error).
Under assumption 3.1, let $\iota>0$ and use Algorithm 3 as instantiations of $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$. If $\mathcal{D}$ is i.i.d. drawn from distribution $\mathcal{P}$, we have:
$(1)$ for any $x_{t}$, we have $\operatorname*{\mathbb{E}}[\mathcal{O}_{1}(x_{t})]=\nabla F(x_{t})$ and
$\Pr[\|\mathcal{O}_{1}(x_{t})-\nabla F(x_{t})\|\geq\zeta_{1}]\leq\iota,$
where $\zeta_{1}=O((G\sqrt{\log(d/\iota)/b})$.
$(2)$ for any $x_{t},x_{t-1}$, we have $\operatorname*{\mathbb{E}}[\mathcal{O}_{2}(x_{t},x_{t-1})]=\nabla F(x_{t})-\nabla F(x_{t-1})$ and
$\Pr[\|\mathcal{O}_{2}(x_{t},x_{t-1})-(\nabla F(x_{t})-\nabla F(x_{t-1}))\|\geq\zeta_{2}]\leq\iota,$
where $\zeta_{2}=O(M\|x_{t}-x_{t-1}\|\sqrt{\log(d/\iota)/b_{t}})$.
Proof.
For each data $z\sim\mathcal{P}$, we know
$\operatorname*{\mathbb{E}}\nabla f(x_{t};z)-\nabla F(x_{t})=0,\quad\|\nabla f(x_{t};z)-\nabla F(x_{t})\|\leq 2G.$
Then the conclusion follows from Lemma 2.4.
Similarly, for each data $z\sim\mathcal{P}$, we know
$\operatorname*{\mathbb{E}}(\nabla f(x_{t};z)-\nabla f(x_{t-1};z))-(\nabla F(x_{t};z)-\nabla F(x_{t-1};z))=0,$
$\|(\nabla f(x_{t};z)-\nabla f(x_{t-1};z))-(\nabla F(x_{t};z)-\nabla F(x_{t-1};z))\|\leq 2M\|x_{t}-x_{t-1}\|.$
The statement (2) also follows from Lemma 2.4. ∎
From now on, we adopt Algorithm 3 as the gradient oracles for Line 5 and Line 7 respectively in Algorithm 2, and we set $\eta=1/M$. We then bound the error between gradient estimator $\nabla_{t}$ and the true gradient $\nabla F(x_{t})$ for Algorithm 2.
Lemma 3.6.
Suppose the dataset is drawn i.i.d. from the distribution $\mathcal{P}$. For any $1\leq t\leq T$ and letting $\tau_{t}\leq t$ be the largest integer such that $\mathrm{drift}_{\tau_{t}}$ is set to be 0, with probability at least $1-T\iota$, for some universal constant $C>0$, we have
$\|\nabla_{t}-\nabla F(x_{t})\|^{2}\leq C(\frac{G^{2}}{b}+\sum_{i=\tau_{t}+1}^{t}M^{2}\|x_{i}-x_{i-1}\|^{2}/b_{i})\log(Td/\iota).$ (1)
Proof.
We consider the case when $t=\tau_{t}$ first, i.e., we query $\mathcal{O}_{1}$ to get $\nabla_{t}$. Then Equation (1) follows from Lemma 3.5.
When $t>\tau_{t}$, then for each $i$ such that $\tau_{t}<i\leq t$, we know conditional on $\nabla_{i-1}$, we have
$\operatorname*{\mathbb{E}}[\Delta_{i}\mid\nabla_{i-1}]=\nabla F(x_{i})-\nabla F(x_{i-1}).$
That is $\Delta_{i}-(\nabla F(x_{i})-\nabla F(x_{i-1}))$ is zero-mean and $\mathrm{nSG}(M\|x_{i}-x_{i-1}\|\sqrt{\log(dT/\iota)}/\sqrt{b_{i}})$ by applying Lemma 3.5. Then Equation (1) follows from applying Lemma 2.4. ∎
Now, we consider the noise added from the tree mechanism to make the algorithm private.
Lemma 3.7.
If we set $\sigma=(\frac{G}{b}+\max_{t}\frac{\|\widetilde{\nabla}_{t}\|}{b_{t}})\log(1/\delta)/\varepsilon$, in the tree mechanism (Algorithm 1) and use Algorithm 3 as gradient oracles, then Algorithm 2 is $(\varepsilon,\delta)$-DP.
Proof.
It suffices to consider the sensitivity of the gradient oracles.
Consider the sensitivity of $\mathcal{O}_{1}$ first. Let $\mathcal{O}(x_{t})^{\prime}$ denote the output with the neighboring dataset. Then it is obvious that
$\|\mathcal{O}_{1}(x_{t})-\mathcal{O}_{1}(x_{t})^{\prime}\|\leq\frac{G}{b}.$
As for the sensitivity of $\mathcal{O}_{2}$, we have
$\|\mathcal{O}_{2}(x_{t},x_{t-1})-\mathcal{O}_{2}(x_{t},x_{t-1})^{\prime}\|\leq\frac{M\|x_{t}-x_{t-1}\|}{b_{t}}=\frac{\|\widetilde{\nabla}_{t}\|}{b_{t}}.$
The privacy guarantee follows from the tree mechanism (Theorem 2.2). ∎
With the noise added from the tree mechanism in mind, now we get the high-probability error bound of our gradient estimators ~$\widetilde{\nabla}_{t}$.
Lemma 3.8.
In Algorithm 2, setting $b=G\sqrt{d}/\varepsilon\alpha+G^{2}/\alpha^{2},b_{t}=\max\{\frac{\|\widetilde{\nabla}_{t}\|\cdot\sqrt{d}}{\alpha\varepsilon},\frac{\kappa\cdot\|\widetilde{\nabla}_{t}\|}{\alpha^{2}},1\}$, and $\sigma=2\log(1/\delta)\alpha/\sqrt{d}$ correspondingly according to Lemma 3.7, for each $t\in[T]$, we know Algorithm 2 is $(\varepsilon,\delta)$-DP, and with probability at least $1-\iota$, ~$\|\widetilde{\nabla}_{t}-\nabla F(x_{t})\|\leq\gamma$, where $\gamma=\tilde{O}(\alpha).$
Proof.
By our setting of parameters, we know
$(\frac{G}{b}+\max_{t}\frac{\|\widetilde{\nabla}_{t}\|}{b_{t}})\leq 2\varepsilon\alpha/\sqrt{d}.$
Then our choice of $\sigma$ ensures the privacy guarantee by Lemma 3.7.
For any $t\in[T]$, we have
~$\|\widetilde{\nabla}_{t}-\nabla F(x_{t})\|\leq\underbrace{\|\widetilde{\nabla}_{t}-\nabla_{t}\|}_{(1)}+\underbrace{\|\nabla_{t}-\nabla F(x_{t})\|}_{(2)}.$
By Theorem 2.2, we know
$(1)\leq\max_{t}\|\mathrm{TREE}(t)\|\leq\sigma\sqrt{d\log(T)}\leq\sigma\sqrt{d\log n}\lesssim\alpha\sqrt{\log n}\leq\tilde{O}(\gamma).$
By Lemma 3.6 and our parameter settings, we have
$\|\nabla_{t}-\nabla F(x_{t})\|^{2}\lesssim$ $(\alpha^{2}+\sum_{i=\tau_{t}+1}^{t}\|\widetilde{\nabla}_{t}\|^{2}/b_{t})\log(nd/\iota)$
$\lesssim$ $(\alpha^{2}+\sum_{i=\tau_{t}+1}^{t}\|\widetilde{\nabla}_{t}\|\cdot\min\{\frac{\alpha\varepsilon}{\sqrt{d}},\alpha^{2}/\kappa\})\log(nd/\iota)$
$\leq$ $(\alpha^{2}+\kappa\cdot\min\{\frac{\alpha\varepsilon}{\sqrt{d}},\alpha^{2}/\kappa\})\log(nd/\iota)$
$\lesssim$ $\alpha^{2}\log(nd/\iota).$
Hence we conclude that $(2)\lesssim\alpha\sqrt{\log(nd/\iota)}$, which completes the proof. ∎
We need to show that we can find an $\gamma$-SOSP before we use up all the functions. We need the following technical Lemma:
Lemma 3.9.
Consider the implementation of Algorithm 2. Suppose the size of dataset $\mathcal{D}$ can be arbitrarily large with functions drawn i.i.d. from $\mathcal{P}$, and we run the algorithm until finding an $\tilde{O}(\gamma)$-SOSP, then with probability at least $1-T\iota$, the total number of functions we will use is bounded by
$\tilde{O}\Big{(}\frac{bBM}{\kappa\gamma}+BM(\frac{\sqrt{d}}{\gamma^{2}\varepsilon}+\frac{\kappa}{\gamma^{3}}+\frac{1}{\gamma^{2}})\Big{)}.$
Proof.
By Proposition 3.4, setting $T=\tilde{O}(B/\eta\gamma^{2})$ suffices to find an $\tilde{O}(\gamma)$-SOSP. Let $\{x_{1},\cdots,x_{t}\}$ be the outputs of the algorithms, where $t\leq T$ denotes the step we halt the algorithm. We show
$\sum_{i=1}^{t}\|\widetilde{\nabla}_{i}\|_{2}\lesssim\tilde{O}(BM/\gamma).$ (2)
Denote the set $S:=\{i\in[t]:\|\widetilde{\nabla}_{i}\|\leq\gamma\}$. As $|S|\leq T=\tilde{O}(B/\eta\gamma^{2})$, we know $\sum_{i\in S}\|\widetilde{\nabla}_{i}\|\leq\tilde{O}(B/\eta\gamma)$.
Now consider the set $S^{c}:=[t]\setminus S$ denoting the index of steps when the norm of the gradient estimator is large. It suffices to bound $\sum_{i\in S^{c}}\|\widetilde{\nabla}_{i}\|_{2}$.
By Lemma 3.2, we know when ~$\|\widetilde{\nabla}_{i}\|\geq\gamma$, $F(x_{i+1})-F(x_{i})\leq-\eta\|\widetilde{\nabla}_{i}\|^{2}/16$, and when ~$\|\widetilde{\nabla}_{i}\|\leq\gamma,F(x_{i+1})\leq F(x_{i})+\eta\gamma^{2}$. Given the bound on the function values, we know
$\sum_{i\in S^{c}}\|\widetilde{\nabla}_{i}\|^{2}\leq\tilde{O}(B/\eta).$
Hence
$\sum_{i\in S^{c}}\|\widetilde{\nabla}_{i}\|\leq\frac{\sum_{i\in S^{c}}\|\widetilde{\nabla}_{i}\|^{2}}{\gamma}\leq\tilde{O}(\frac{B}{\eta\gamma}).$
This completes the proof of Equation (2).
The total number of functions we used for $\mathcal{O}_{1}$ is upper bounded by $b\cdot\frac{\sum_{i\in[t]\|\widetilde{\nabla}_{i}\|}}{\kappa}=\tilde{O}(bBM/\kappa\gamma)$. The total number of functions we used for $\mathcal{O}_{2}$ is upper bounded as follows:
$\sum_{i\in[t]}b_{t}\lesssim(\frac{\sqrt{d}}{\alpha\varepsilon}+\frac{\kappa}{\alpha^{2}})\sum_{i\in[t]}\|\widetilde{\nabla}_{t}\|+T\leq BM\cdot\tilde{O}(\frac{\sqrt{d}}{\gamma^{2}\varepsilon}+\frac{\kappa}{\gamma^{3}}+\frac{1}{\gamma^{2}}).$
This completes the proof. ∎
Given the dataset size requirement, we can get the final bound on finding SOSP.
Lemma 3.10.
With $\mathcal{D}$ of size $n$ drawn i.i.d. from $\mathcal{P}$, setting $\kappa=\max\{\frac{\alpha\sqrt{d}}{\varepsilon},(BGM)^{1/3}\}$,
$\alpha=O\Big{(}((BGM)^{1/3}+\sqrt{BM})(\frac{\sqrt{d}}{n\varepsilon})^{1/2}+\frac{B^{\frac{2}{9}}M^{\frac{2}{9}}G^{\frac{5}{9}}+B^{\frac{4}{9}}M^{\frac{4}{9}}G^{\frac{1}{9}}}{n^{1/3}}+\frac{(MB)^{1/2}}{\sqrt{n}}\Big{)},$
and other parameters as in Lemma 3.8, with probability at least $1-\iota$, at least one of the outputs of Algorithm 2 is $\gamma$-SOSP, with $\gamma=\tilde{O}(\alpha)$.
Proof.
By Lemma 3.9 and our parameter settings, we need
$n\geq\tilde{\Omega}\big{(}\frac{bBM}{\kappa\gamma}+BM(\frac{\sqrt{d}}{\gamma^{2}\varepsilon}+\frac{\kappa}{\gamma^{3}}+\frac{1}{\gamma^{2}})\big{)}.$
First,
$n\geq\tilde{\Omega}(\frac{bBM}{\kappa\gamma})=\Theta(\frac{BGM\sqrt{d}}{\kappa\varepsilon\gamma^{2}}+\frac{G^{2}BM}{\kappa\gamma^{3}})\Leftarrow\gamma\geq\tilde{O}(\frac{(BGM)^{1/3}d^{1/4}}{\sqrt{n\varepsilon}}+\frac{B^{\frac{2}{9}}M^{\frac{2}{9}}G^{\frac{5}{9}}}{n^{1/3}}).$
Secondly,
$n\geq\tilde{\Omega}(BM\sqrt{d}/\gamma^{2}\varepsilon)\Leftarrow\gamma\geq\tilde{O}(\sqrt{BM}(\frac{\sqrt{d}}{n\varepsilon})^{1/2}).$
Thirdly,
$n\geq\tilde{\Omega}(BM\kappa/\gamma^{3})\Leftrightarrow n\geq\tilde{\Omega}(BM(\frac{\sqrt{d}}{\gamma^{2}\varepsilon})+B^{4/3}M^{4/3}G^{1/3}/\gamma^{3})$
$\Leftarrow\gamma\geq\tilde{O}(\sqrt{BM}(\frac{\sqrt{d}}{n\varepsilon})^{1/2}+\frac{B^{\frac{4}{9}}M^{\frac{4}{9}}G^{\frac{1}{9}}}{n^{1/3}}).$
Finally,
$n\geq\tilde{\Omega}(MB/\gamma^{2})\Leftarrow\gamma\geq\tilde{O}(\frac{(MB)^{1/2}}{\sqrt{n}}).$
Combining these together, we get the claimed statement.
Combining Lemma 3.7 and Lemma 3.10, we have the following main result for finding SOSP privately:
Theorem 3.11.
Given $\varepsilon,\delta>0$, with gradient oracles in Algorithm 3, setting $b=G(\sqrt{d}/\varepsilon\alpha+1/\alpha^{2}),b_{t}=\max\{\frac{\|\widetilde{\nabla}_{t}\|\cdot\sqrt{d}}{\alpha\varepsilon},\frac{\|\widetilde{\nabla}_{t}\|\kappa}{\alpha^{2}},1\}$, $\kappa=\max\{\frac{\alpha\sqrt{d}}{\varepsilon},(BGM)^{1/3}\}$ and $\sigma=\alpha/\sqrt{d}$, Algorithm 2 is $(\varepsilon,\delta)$-DP, and if the dataset is i.i.d. drawn from the underlying distribution $\mathcal{P}$, at least one of the output of is $\tilde{O}(\alpha)$-SOSP, where
$\alpha=O\Big{(}((BGM)^{1/3}+\sqrt{BM})(\frac{\sqrt{d}}{n\varepsilon})^{1/2}+\frac{B^{\frac{2}{9}}M^{\frac{2}{9}}G^{\frac{5}{9}}+B^{\frac{4}{9}}M^{\frac{4}{9}}G^{\frac{1}{9}}}{n^{1/3}}+\frac{(MB)^{1/2}}{\sqrt{n}}\Big{)}$
Remark 3.12.
If we treat the parameters $B,G,M$ as constants $O(1)$, then we get $\alpha=\tilde{O}((\frac{\sqrt{d}}{n\varepsilon})^{1/2}+\frac{1}{n^{1/3}})$ as claimed before in the abstract and introduction.
If we make further assumptions, like assuming the functions are defined over a constraint domain $\mathcal{X}\subset\mathbb{R}^{d}$ of diameter $R$ and we allow exponential running time, we can get some other standard bounds that can be better than Theorem 3.11 in some regimes. See Appendix A.2 for more discussions.
4 Discussion
We combine the concepts of adaptive batch sizes and the tree mechanism to improve the previous best results for privately finding SOSP. Our approach achieves the same bound as the state-of-the-art method for finding FOSP, suggesting that privately finding an SOSP may incur no additional cost.
Several interesting questions remain. First, what is the tight bound for privately finding FOSP and SOSP? Second, can the adaptive batch size technique be applied in other settings? Could it offer additional advantages, such as reducing runtime in practice? Finally, while we can obtain a generalization error bound of $\sqrt{d/n}$ using concentration inequalities and the union bound, can we achieve a better generalization error bound for the non-convex optimization?
Appendix A Appendix
A.1 Discussion of Proposition 3.4
Proposition 3.4 has become a standard result in non-convex optimization. Ganesh et al. [2023] assume two kinds of gradient oracles that satisfy the norm-Subgaussian definition below:
Definition A.1 (SubGaussian gradient oracles).
For a $G$-Lipschitz and $M$-smooth function $F$:
$(1)$ We say $\mathcal{O}_{1}$ is a first kind of $\zeta_{1}$ norm-subGaussian Gradient oracle if given $x\in\mathbb{R}^{d}$, $\mathcal{O}(x)$ satisfies $\operatorname*{\mathbb{E}}\mathcal{O}_{1}(x)=\nabla F(x)$ and $\mathcal{O}_{1}(x)-\nabla F(x)$ is $\mathrm{nSG}(\zeta_{1})$.
$(2)$ We say $\mathcal{O}_{2}$ is a second kind of $\zeta_{2}$ norm-subGaussian stochastic Gradient oracle if given $x,y\in\mathbb{R}^{d}$, $\mathcal{O}_{2}(x,y)$ satisfies that $\operatorname*{\mathbb{E}}\mathcal{O}_{2}(x,y)=\nabla F(x)-\nabla F(y)$ and $\mathcal{O}_{2}(x,y)-(\nabla F(x)-\nabla F(y))$ is $\mathrm{nSG}(\zeta_{2}\|x-y\|)$.
In Definition A.1, the oracles are required to be unbiased with norm-subGaussian error. While this condition is sufficient, it is not necessary. These two types of gradient oracles are then used to construct gradient estimators whose errors are tightly controlled with high probability (Lemma3.3 in Ganesh et al. [2023]). Proposition 3.4 can be established directly using Lemma 3.3 from Ganesh et al. [2023], as demonstrated in the proof of Lemma 3.6 in the same work. We include the proof here for completeness.
Proof of Proposition 3.4.
By Lemma 3.2 and the precondition that ~$\|\widetilde{\nabla}_{t}-\nabla F(x_{t})\|\leq\gamma$, we know that, if $\|\nabla F(x_{t})\|\geq 4\gamma$, then $F(x_{t+1})-F(x_{t})\leq-\eta\|\widetilde{\nabla}\|^{2}/16$. Otherwise, $\|\nabla F(x_{t})\|<4\gamma$. If $\|\nabla F(x_{t})\|<4\gamma$ but $x_{t}$ is a saddle point, then by Lemma 3.3, we know
$F(x_{t+\Gamma})-F(x_{t})\leq-\tilde{\Omega}(\gamma^{3/2}/\sqrt{\rho}),$
where $\Gamma=\tilde{O}(M/\sqrt{\rho\gamma})$. Then if none of point in $\{x_{i}\}_{i\in[T]}$ is an $\tilde{O}(\gamma)$-SOSP, then we know $F(x_{T})-F(x_{0})<-B$, which is contradictory to Assumption 3.1. Hence at least one point in $\{x_{i}\}_{i\in[T]}$ should be an $\tilde{O}(\gamma)$-SOSP, and hence complete the proof.
A.2 Other results
The first result is combining the current result in finding the SOSP of the empirical function $F_{\mathcal{D}}(x):=\frac{1}{n}\sum_{\zeta\in\mathcal{D}}f(x;\zeta)$, and then apply the generalization error bound as follows:
Theorem A.2.
Suppose $\mathcal{D}$ is i.i.d. drawn from the underlying distribution $\mathcal{P}$ and under Assumption 3.1. Additionaly assume $f(;\zeta):\mathcal{X}\to\mathbb{R}$ for some constrained domain $\mathcal{X}\subset\mathbb{R}^{d}$ of diameter $R$. Then we know for any point $x\in\mathcal{X}$, with probability at least $1-\iota$, we have
$\|\nabla F_{\mathcal{P}}(x)-\nabla F_{\mathcal{D}}(x)\|\leq\tilde{O}(\sqrt{d/n}),\|\nabla^{2}F_{\mathcal{P}}(x)-\nabla^{2}F_{\mathcal{D}}(x)\|\leq\tilde{O}(\sqrt{d/n}).$
Proof.
We construct a maximal packing $\mathcal{Y}$ of $O((R/r)^{d})$ points for $\mathcal{X}$, such that for any $x\in\mathcal{X}$, there exists a point $y\in\mathcal{Y}$ such that $\|x-y\|\leq r$.
By Union bound, the Hoeffding inequality for norm-subGaussian (Lemma 2.4 and the Matrix Bernstein Inequality(Theorem 2.5), we know with probability at least $1-\tau$, for all point $y\in\mathcal{Y}$, we have
$\|\nabla F_{\mathcal{P}}(y)-\nabla F_{\mathcal{D}}(y)\|\leq\tilde{O}(L\sqrt{d\log(R/r)/n}),\|\nabla^{2}F_{\mathcal{P}}(y)-\nabla^{2}F_{\mathcal{D}}(y)\|\leq\tilde{O}(M\sqrt{d\log(R/r)/n}).$ (3)
Conditional on the above event Equation (3). Choosing $r\leq\min\{1,M/\rho\}\sqrt{d/n}$, then by the assumptions on Lipschitz and smoothness, we have for any $x\in\mathcal{X}$, there exists $y\in\mathcal{Y}$ such that $\|x-y\|\leq r$, and
$\|\nabla F_{\mathcal{P}}(x)-\nabla F_{\mathcal{D}}(x)\|\leq$ $\|\nabla F_{\mathcal{P}}(x)-\nabla F_{\mathcal{P}}(y)\|+\|\nabla F_{\mathcal{P}}(y)-\nabla F_{\mathcal{D}}(y)\|+\|\nabla F_{\mathcal{D}}(y)-\nabla F_{\mathcal{D}}(x)\|$
$\leq$ $\tilde{O}(L\sqrt{d/n}).$
Similarly, we can show
$\|\nabla^{2}F_{\mathcal{P}}(x)-\nabla^{2}F_{\mathcal{D}}(x)\|\leq\tilde{O}((M+\rho r)\sqrt{d/n})=\tilde{O}(M\sqrt{d/n}).$
The current SOTA of finding SOSP privately of $F_{\mathcal{D}}$ is from Ganesh et al. [2023], where they can find an $\tilde{O}((\sqrt{d}/n\varepsilon)^{2/3})$-SOSP. Combining the SOTA and Theorem A.2, we can find the $\alpha$-SOSP of $F_{\mathcal{P}}$ privately with
$\alpha=\tilde{O}(\frac{\sqrt{d}}{n}+(\frac{\sqrt{d}}{n\varepsilon})^{2/3}).$
If we allow exponential running time, as Lowy et al. [2024] suggests, we can find an initial point $x_{0}$ privately to minimize the empirical function and then use $x_{0}$ as a warm start to improve the final bound further.